Parth (parthkohli):

Just like there are imaginary numbers, are there some famous imaginary functions too?

OpenStudy (fellowroot):



OpenStudy (anonymous):

theres ones relating e, i, and sin/cos... more info, you can start here: http://en.wikipedia.org/wiki/Complex_analysis

hartnn (hartnn):

isn't \(\large f(x)=e^{i\theta}\) pretty famous?

OpenStudy (anonymous):

also, e^(i(pi))-1=0

Parth (parthkohli):

That's not an imaginary function. It's legit.

Parth (parthkohli):

@etemplin e^(i * pi) + 1 = 0

Parth (parthkohli):

Like a function which doesn't exist, but is assumed to be.

OpenStudy (anonymous):

oh sorry...i was close

Parth (parthkohli):

Wait, that actually doesn't exist. You're right. :D

Parth (parthkohli):

Hehe, sorry.

Parth (parthkohli):

And thanks!